OpenStudy (anonymous):

Can anyone help me get started or give me a hint on tan(arcsin(3/8)). Thanks in advance!!

OpenStudy (freckles):

let y=arcsin(3/8) so then sin(y)=3/8 draw a right triangle and use sin(y)=3/8 and then find the adjacent side tan(y)=opp/adj